\begin{thebibliography}{99}

\bibitem{bamler2016proof}
R. H. Bamler,
A Ricci flow proof of a result by Gromov on lower bounds for scalar curvature,
Math. Res. Letters \textbf{23}, 325--337 (2016).


\bibitem{besson1996riemannian}
G. Besson, J. Lohkamp, P. Pansu and P. Petersen,
Riemannian geometry,
Fields institute monographs vol. \textbf{4},
American Mathematical Society (1996).


\bibitem{brendle2005convergence}
S. Brendle,
Convergence of the Yamabe flow for arbitrary initial energy,
J. Differential Geom. \textbf{69}, 217--278 (2005).



\bibitem{caldeira2021normalized}
B. Caldeira, L. Hartmann and B. Vertman,
Normalized Yamabe flow on manifolds with bounded geometry,
Ann. Glob. Anal. Geom. \textbf{63}, 23 (2023).

\bibitem{carron2021convergence}
G. Carron, J. O. Lye and B. Vertman,
Convergence of the Yamabe flow on singular spaces with positive Yamabe constant,
Tohoku Math. J. \textbf{75}, 561--615 (2023).

\bibitem{chow1992yamabe}
B. Chow,
The Yamabe flow on locally conformally flat manifolds with positive Ricci curvature,
Comm. Pure Appl. Math. \textbf{45}, 1003--1014 (1992).


\bibitem{gromov2014dirac}
M. Gromov,
Dirac and Plateau billiards in domains with corners,
Cent. Eur. J. Math. \textbf{12}, 1109--1156 (2014).


%\bibitem{hamanaka2022c}
%S. Hamanaka,
%Limit theorems for the total scalar curvature,
%arXiv preprint arXiv:2208.01865v14 (2022).
%(old version: $C^{0}, C^{1}$-limit theorems for total scalar curvatures)

%\bibitem{hamilton1989lectures}
%R. S. Hamilton,
%Lectures on geometric flows, 
%unpublished manuscript, 1989.

\bibitem{krylov1981certain}
N. V. Krylov and M. V. Safonov,
A certain property of solutions of parabolic equations with measurable coefficients,
Mathematics of the USSR-Izvestiya \textbf{16}, 151--164 (1981).



\bibitem{lohkamp1995curvature}
J. Lohkamp,
Curvature h-principles,
Ann. of Math. \textbf{142}, 457--498 (1995).

\bibitem{picard2019notes}
S. Picard,
Notes on H{\"o}lder Estimates for Parabolic PDE lecture notes,
available online in \url{https://people.math.harvard.edu/~spicard/notes-parabolicpde.pdf}.


\bibitem{schwetlick2003convergence}
H. Schwetlick and M. Struwe,
Convergence of the Yamabe flow for ``large'' energies,
J. Reine Angew. Math. \textbf{562}, 59--100 (2003).

\bibitem{takahashi2022infinite}
J. Takahashi and H. Yamamoto,
Infinite-time incompleteness of noncompact Yamabe flow,
Calc. Var. Partial Differential Equations \textbf{61}, 1--24 (2022).


\bibitem{ye1994global}
R. Ye,
Global existence and convergence of Yamabe flow,
J. Differential Geom. \textbf{39}, 35--50 (1994).
\end{thebibliography}

\documentclass[10pt]{article}

\usepackage{bigints}

\usepackage{amssymb}
\usepackage{mathrsfs}
\usepackage{amsmath}
\usepackage{amsfonts}

\renewcommand{\thefootnote}{\fnsymbol{footnote}}

\usepackage{amsthm}
\newtheorem{theo}{Theorem}[section]
\newtheorem{prop}{Proposition}[section]
\newtheorem{coro}{Corollary}[section]
\newtheorem{lemm}{Lemma}[section]

\newtheorem{main}{Main Theorem}

\theoremstyle{definition}

\theoremstyle{remark}

\newtheorem{rema}{Remark}[section]
\newtheorem{ques}{Question}[section]

\usepackage{hyperref}

\title{Upper bound preservation of the total scalar curvature in a conformal class}
\author{Shota Hamanaka\thanks{The author was supported by Foundation of Research Fellows, The Mathematical Society of Japan.}}
\date{\today}

\begin{document}
\maketitle
 
\renewcommand{\thefootnote}{\fnsymbol{footnote}} 
\footnotetext{\emph{Keywords}: Scalar curvature, Yamabe flow}     
\renewcommand{\thefootnote}{\arabic{footnote}}

\renewcommand{\thefootnote}{\fnsymbol{footnote}} 
\footnotetext{\emph{2020 Mathematics Subject Classification}: 53C21, 53E20}     
\renewcommand{\thefootnote}{\arabic{footnote}}
  
\begin{abstract}
We show that in an arbitrarily fixed conformal class on a closed manifold, the upper bound condition of the total scalar curvature is $C^{0}$-closed if its Yamabe constant is non-positive.
Moreover, if we consider the condition that the scalar curvature is bounded by some fixed continuous function from below in addition to the upper bound of the total scalar curvature, then such a condition is $C^{0}$-closed in an arbitrarily fixed conformal class, whose Yamabe constant is positive.
\end{abstract}

\section{Introduction}
\,\,\,\,\,\,\, Let $M^{n}$ be a smooth manifold of dimension $n \ge 2$ and $\mathcal{M}$ the space of all $C^{2}$-Riemannian metrics on $M.$
For example, the followings are known for the topology of several subspaces of $\mathcal{M}$.
\begin{itemize}
\item Gromov \cite{gromov2014dirac} and Bamler \cite{bamler2016proof} proved that for any continuous function $\sigma \in C^{0}(M),$ the space
\[
\left\{ g \in \mathcal{M}~\middle|~R(g) \ge \sigma \right\}
\]
is closed in $\mathcal{M}$ with respect to the $C^{0}$-topology,
i.e., if a sequence of $C^{2}$-metrics $g_{i}$ converges to a $C^{2}$-metric $g$ in the local $C^{0}$-sense on $M$ that satisfy $R(g_{i}) \ge \sigma,$
then $R(g) \ge \sigma.$
Here, $R(g)$ denotes the scalar curvature of $g$.
\item Lohkamp \cite{lohkamp1995curvature} proved that for any continuous function $\sigma \in C^{0}(M),$ the space
\[
\mathcal{R}(M)_{\le \sigma} := \left\{ g \in \mathcal{M}~\middle|~R(g) \le \sigma \right\}
\]
is dense in $\mathcal{M}$ with respect to the $C^{0}$-topology,
i.e., for any $C^{2}$-metric $g,$ there exists a sequence of $C^{2}$-metrics $(g_{i})$ satisfying $R(g_{i}) \le \sigma$ such that $g_{i}$ converges to $g$ in the locally $C^{0}$-sense on $M$.

This result also implies that, when $M$ is closed (i.e., compact without boundary), for any constant $\kappa \in \mathbb{R},$ the space
\[
\left\{ g \in \mathcal{M}~\middle|~\int_{M} R(g)\, d\mathrm{vol}_{g} \le \kappa \right\}
\]
is dense in $\mathcal{M}$ with respect to the $C^{0}$-topology.
Here, $d\mathrm{vol}_{g}$ denotes the Riemannian volume measure of $g.$
\begin{proof}
  From the above fact by Lohkamp, we can take a sequence of metrics $(g_{i})$ such that
\[
R(g_{i}) \le \frac{\kappa}{2 \mathrm{Vol}(M, g)}~~\mathrm{if}~\kappa \ge 0~~\left(\mathrm{resp.}~R(g_{i}) \le \frac{2 \kappa}{\mathrm{Vol}(M, g)}~~\mathrm{if}~\kappa < 0 \right),
\]
and $g_{i}$ converges to $g$ in the $C^{0}$-sense on $M.$
Integrating both sides with respect to $g_{i}$, we get
\[
\int_{M} R(g_{i})\, d\mathrm{vol}_{g_{i}} \le \frac{\mathrm{Vol}(M, g_{i})}{2 \mathrm{Vol}(M, g)} \kappa~~\mathrm{if}~\kappa \ge 0~~\left( \mathrm{resp.}~\le \frac{2 \mathrm{Vol}(M, g_{i})}{\mathrm{Vol}(M, g)} \kappa~~\mathrm{if}~\kappa < 0 \right),
\]
Since $g \mapsto \mathrm{Vol}(M, g)$ is continuous with respect to $C^{0}$-topology, hence $\mathrm{Vol}(M, g_{i}) / \mathrm{Vol}(M, g) \rightarrow 1$ as $i \rightarrow \infty.$
Therefore there is a sufficiently large $i_{0} \in \mathbb{N}$ such that for all $i \ge i_{0}$, $\int_{M} R(g_{i})\, d\mathrm{vol}_{g_{i}} \le \kappa$, and hence $(g_{i})_{i \ge i_{0}}$ is the desired $C^{0}$-approximation of $g$.
\end{proof}
This Lohkamp's approximation was built from certain $C^{0}$-deformation which is not just a conformal deformation.
Indeed, in the present paper, we will show that in a fixed conformal class, such a subspace is $C^{0}$-closed in some sense (see Corollary \ref{coro1} below).
\item Lohkamp \cite{besson1996riemannian} proved that for any continuous function $\sigma \in C^{0}(M),$ both of the spaces
\[
\left\{ g \in \mathcal{M}~\middle|~\mathrm{Ric}(g) \le \sigma \cdot g \right\}
\]
and
\[
\left\{ g \in \mathcal{M}~\middle|~R(g) \le \sigma \right\}
\]
are closed in $\mathcal{M}$ with respect to $C^{1}$-topology.
\item The author (in his preprint, arXiv:2208.01865v14) proved that when $M$ is closed, for any continuous nonnegative function $\sigma \in C^{0}(M, \mathbb{R}_{\ge 0})$ and constant $\kappa \in \mathbb{R},$ the space
\[
\left\{ g \in \mathcal{M}~\middle|~\int_{M} R(g)\, d\mathrm{vol}_{g} \ge \kappa,~R(g) \ge \sigma \right\}
\]
is closed in $\mathcal{M}$ with respect to the $W^{1, p}~(p > n)$-topology.
\end{itemize}
In the present paper, we will examine the space of all Riemannian metrics whose total scalar curvatures are bounded from \textbf{above} by some constant.
Our first main result is the following.
\begin{main}
\label{main}
Let $M^{n}$ be a closed manifold of dimension $n \ge 3$ and $g_{0}$ a $C^{2}$-Riemannian metric on $M.$
Let $u_{i}, u : M \rightarrow \mathbb{R}_{+}~(i = 1,2, \cdots )$ be positive $C^{2}$-functions on $M$
and assume the following: 
\begin{itemize}
\item[$(1)$] $g_{i} := u_{i}^{\frac{4}{n-2}} g_{0} \overset{C^{0}}{\longrightarrow} g := u^{\frac{4}{n-2}} g_{0}$ on $M,$
\item[$(2)$] there is a constant $\kappa \in \mathbb{R}$ such that $\int_{M} R(g_{i})\, d\mathrm{vol}_{g_{i}} \le \kappa~~\mathrm{for~all}~i.$
\end{itemize}
When $(M, g_{0})$ is Yamabe positive, i.e., $Y(M, g_{0}) > 0$ (see Section \ref{section-proof} for the definition of the Yamabe constant $Y(M, g_{0})$), we additionally assume that 
\begin{itemize}
\item[$(3)$] there is a continuous function  $\delta \in C^{0}(M)$ such that for all $i$, $R(g_{i}) \ge \delta$ on $M.$
\end{itemize}
Then $\int_{M} R(g)\, d\mathrm{vol}_{g} \le \kappa.$
Here, $R(g),$ $d\mathrm{vol}_{g}$ denote respectively the scalar curvature and the Riemannian volume measure of $g.$
Moreover, when $(M, g_{0})$ is Yamabe positive, the same assertion still holds even if $(1)$ is replaced with the following weaker condition:
\begin{itemize}
\item[$(1)'$] $u_{i} \rightarrow u$ in the $L^{\frac{2n}{n-2}}(M, g_{0})$-sense, i.e., 
\[
\int_{M} \left| u_{i} - u \right|^{\frac{2n}{n-2}} d\mathrm{vol}_{g_{0}} \rightarrow 0~~~\mathrm{as}~i \rightarrow \infty.
\]
\end{itemize}
\end{main}
As a corollary of Main Theorem \ref{main} combined with {\cite[p.1118]{gromov2014dirac}}, we obtain the following.
\begin{coro}
\label{coro1}
Let $M^{n}$ be a closed manifold of dimension $n \ge 2$ and $g_{0}$ a $C^{2}$-Riemannian metric on $M.$
Let $\mathcal{M}$ be the space of all $C^{2}$-Riemannian metrics on $M.$
Then, for any continuous function $\sigma \in C^{0}(M)$ and constant $\kappa \in \mathbb{R},$ the space
\[
\left\{ g \in [g_{0}]~\middle|~\int_{M} R(g)\, d\mathrm{vol}_{g} \le \kappa,~R(g) \ge \sigma \right\}~\left( \subset \mathcal{M} \right)
\]
when $Y(M, g_{0}) > 0,$ and the space 
\[
\left\{ g \in [g_{0}]~\middle|~\int_{M} R(g)\, d\mathrm{vol}_{g} \le \kappa \right\}~\left( \subset \mathcal{M} \right)
\]
when $Y(M, g_{0}) \le 0$ are both closed in $\mathcal{M}$ with respect to $C^{0}$-topology.
Here, $[g_{0}] := \{ g = u \cdot g_{0} \in \mathcal{M}~|~u \in C^{2}(M),~u > 0~\mathrm{on}~M \}$ is the ($C^{2}$-) conformal class of $g_{0}.$
\end{coro}
When $n = 2,$ the above corollary is a direct conclusion from the Gauss-Bonnet Theorem and {\cite[p.1118]{gromov2014dirac}}.
In the case of $\kappa \le 0$ (in particular, $Y(M, g_{0}) \le 0$), we can refine Main Theorem \ref{main} to the following one with a slightly weaker assumption. 
\begin{main}
\label{main2}
Let $M^{n}$ be a closed manifold of dimension $n \ge 3$ and $g_{0}$ a $C^{2}$-Riemannian metric on $M.$
Let $u_{i}, u : M \rightarrow \mathbb{R}_{+}~(i = 1,2, \cdots )$ be positive $C^{2}$-functions on $M$
and set $g_{i} := u_{i}^{\frac{4}{n-2}} g_{0}$ and $g := u^{\frac{4}{n-2}} g_{0}.$
Assume the following:
\begin{itemize}
\item[$(a)$] there is a positive constant $C_{0}$ such that $C_{0}^{-1} \le u_{i} \le C_{0}$ on $M$ for all $i,$
\item[$(b)$] $u_{i} \rightarrow u$ in the $L^{1}(M, g_{0})$-sense, i.e., 
\[
\int_{M} |u_{i} - u|\, d\mathrm{vol}_{g_{0}} \rightarrow 0~~~\mathrm{as}~i \rightarrow \infty,
\]
\item[$(c)$] there is a non-positive constant $\kappa \in \mathbb{R}$ such that $\int_{M} R(g_{i})\, d\mathrm{vol}_{g_{i}} \le \kappa \le 0$ for all $i$.
In particular, $g_{0}$ is Yamabe non-positive (i.e., $Y(M, g_{0}) \le 0$) from this condition.
\end{itemize}
Then $\int_{M} R(g)\, d\mathrm{vol}_{g} \le \kappa.$
\end{main}

This paper is organized as follows.
In Section 2, we prove Main Theorem \ref{main}. 
The key fact to prove Main Theorem \ref{main} is that the normalized Yamabe flow starting from each $g_{i}$ subconverges to the normalized Yamabe flow starting from the limiting metric $g.$
This has been proven separately in two cases: Yamabe non-positive (i.e., $Y(M, g_{0}) \le 0$) and positive (i.e., $Y(M, g_{0}) > 0$).
In Section 3, we prove Main Theorem \ref{main2}.
In the setting of Main Theorem \ref{main2}, we observe the unnormalized Yamabe flow starting from each $g_{i}.$
We prepare an $L^{1}$-estimate for such flows and use it to prove that $g_{i}$ actually converges to $g$ in the $C^{0}$-sense on $M$ as $i \rightarrow \infty.$
Then Main Theorem \ref{main2} follows from Main Theorem \ref{main}. 
In Section 4 (Appendix), we show an auxiliary fact has been used in the proof of Main Theorem \ref{main}.

\subsection*{Acknowledgements}
\,\,\,\,\,\,\, This work was supported by Foundation of Research Fellows, The Mathematical Society of Japan.
The author was supported by Mitsubishi Electric Corporation Advanced Technology R\&D Center.
The author would like to thank J{\o}rgen Olsen Lye for notifying him of their paper \cite{carron2021convergence}.

\section{Proof of Main Theorem \ref{main}}
\label{section-proof}

\,\,\,\,\,\,\, Let $M$ be a closed (i.e., compact without boundary) manifold of dimension $n \ge 3$
and let $g$ be a Riemannian metric on $M.$
We will denote the Riemannian volume measure of $(M, g)$ as $d\mathrm{vol}_{g}.$
We consider the (volume) normalized Yamabe flow:
\begin{equation}\label{eq-normalized-yamabe}
  \frac{\partial}{\partial t} g(t) = - \left( R(g(t)) - r(g(t)) \right)\, g(t),
\end{equation}
where $R(g(t))$ is the scalar curvature of $g(t)$ and $r(g(t))$ is the mean value of $R(g(t)),$ i.e., 
\[
r(g(t)) = \frac{\int_{M} R(g(t))\, d\mathrm{vol}_{g(t)}}{\int_{M} d\mathrm{vol}_{g(t)}}.
\]
Since 
\[
\begin{split}
\frac{d}{dt} \mathrm{Vol}(M, g(t)) &= \int_{M} \frac{1}{2} \mathrm{tr}_{g(t)} \left( \frac{\partial}{\partial t} g(t) \right)\, d\mathrm{vol}_{g(t)} \\
&= -\frac{n}{2} \int_{M} \left( R(g(t)) - r(g(t)) \right)\, d\mathrm{vol}_{g(t)}
= 0,
\end{split}
\]
the volume $\mathrm{Vol}(M, g(t))$ is invariant along the normalized Yamabe flow.
The Yamabe constant of a Riemannian metric $g_{0}$ is defined as the infimum of the normalized Einstein-Hilbert functional among all metrics conformally equivalent to $g_{0},$ i.e., 
\[
Y(M, g_{0}) := \inf \left\{ \frac{\int_{M} R(g)\, d\mathrm{vol}_{g}}{\left( \int_{M} d\mathrm{vol}_{g} \right)^{\frac{n-2}{n}}}~\middle|~g = u^{\frac{4}{n-2}} g_{0},~u \in C^{\infty}(M),~u > 0~\mathrm{on}~M \right\}.
\]
By the definition, $Y(M, g_{0})$ depends only on the conformal class $[g_{0}]$ of $g_{0}.$

Since the normalized Yamabe flow preserves the conformal structure, we may write $g(t) = u(t)^{\frac{4}{n-2}} g_{0},$ where $g_{0}$ is a fixed background metric on $M$ and $u(t)$ is a positive function.
The scalar curvature of $g(t)$ is related to the scalar curvature of $g_{0}$ by
\[
R(g(t)) = - u^{-\frac{n+2}{n-2}} \left( \frac{4(n-1)}{n-2} \Delta_{g_{0}} u - R(g_{0}) u \right).
\]
Hence, the normalized Yamabe flow reduces to the following evolution equation for the conformal factor can be written as 
\[
\frac{\partial}{\partial t} u(t) = \frac{n+2}{4} \left( \frac{4(n-1)}{n-2}\Delta_{g_{0}}u(t)^{\frac{n-2}{n+2}} - R(g_{0}) u(t)^{\frac{n-2}{n+2}} + r(g(t)) u(t) \right).
\]
Using the identity,
\[
\frac{\partial}{\partial t} R(g(t)) = (n-1) \Delta_{g(t)} R(g(t)) + R(g(t)) \left( R(g(t)) - r(g(t)) \right), 
\]
we obtain that
\[
\frac{d}{dt} r(g(t)) = -\frac{n-2}{2} \mathrm{Vol}(M, g(t))^{-1} \int_{M} \left( R(g(t)) -r(g(t)) \right)^{2}\, d\mathrm{vol}_{g(t)}.
\]
In particular, the function $t \mapsto r(g(t))$ is decreasing.

It is well-known that for a given positive function $u_{0}$ on $M$ with sufficiently high regularity, the normalized Yamabe flow always has a unique positive smooth solution $u(t)$ up to a positive time $T > 0$ with initial data $u_{0}.$
We use the following version by Carron, Lye and Vertman \cite{carron2021convergence}.
\begin{prop}[{\cite[Theorem 2.7, Theorem 2.9]{carron2021convergence}}]
\label{prop-short}
Let $M^{n}$ be a closed manifold of dimension $n \ge 3$ and $p > n/2.$
Let $g_{0}$ be a $C^{2}$-Riemannian metric on $M$ and
let $u_{0}$ be a positive $C^{2}$-function.
Then there is a positive time $T > 0$ such that a unique positive solution $u(t)$ of the normalized Yamabe flow (\ref{eq-normalized-yamabe}) on $M \times [0, T)$, and   the solution $u(t)$ is smooth on $M \times (0, T)$.
Moreover, $u(t) \rightarrow u_{0}$ as $t \rightarrow 0$ in the $C^{0}$-sense
and 
$R(g(t)) = R ( u(t)^{\frac{4}{n-2}} g_{0} )$ converges to $R( u_{0}^{\frac{4}{n-2}} g_{0} )$ as $t \rightarrow 0$
in the $C^{0}$-sense.
\end{prop}
\begin{proof}
Since $(M, g_{0})$ is a closed smooth Riemannian manifold, the assumption of {\cite[Theorem 2.7]{carron2021convergence}} is satisfied.
Hence the assertion follows from {\cite[Theorem 2.9]{carron2021convergence}}.
Note that Assumption 1 in \cite{carron2021convergence} is used to gain the Sobolev inequality.
However, in our setting, the manifold $(M, g_{0})$ is smooth and closed, hence the Sobolev inequality holds regardless of the sign of the Yamabe constant.
As a result, we can obtain the unique short-time solution of (\ref{eq-normalized-yamabe}) for a closed smooth Riemannian manifold $(M, g_{0})$ with $Y(M, g_{0}) \le 0$ as well.
Finally, the last assertion follows from {\cite[The annotation 6 in Theorem 2.9]{carron2021convergence}}.
\end{proof}
\begin{rema}
\label{rema-prop2.1}
Since 
$u(t) \rightarrow u_{0}$ and 
$R(g(t)) \rightarrow R \left( u_{0}^{\frac{4}{n-2}} g_{0} \right)$ in the $C^{0}$-sense as $t \rightarrow 0$, 
from the mean value theorem, we still have that 
$t \mapsto r(g(t))$ is decreasing on $[0, T)$.
\end{rema}
The existence of a long-time solution were settled by Hamilton (in his unpublished manuscript), Chow \cite{chow1992yamabe}, Schwetlick-Struwe \cite{schwetlick2003convergence} and Brendle \cite{brendle2005convergence}.
\begin{theo}[\cite{chow1992yamabe, schwetlick2003convergence, brendle2005convergence},~cf. {\cite[Proof of Proposition 3.6]{carron2021convergence}}]
Let $M, g_{0}$ and $u_{0}$ be the same as those in the previous proposition.
Then the positive solution $u(t)$ as in Proposition \ref{prop-short} exists for all time (i.e., $T = \infty$).
\end{theo}
Therefore, in the setting of Main Theorem \ref{main}, we can take a uniform positive time $T > 0$ such that
the unique positive solutions $u(t), u_{i}(t)$ of the normalized Yamabe flow (\ref{eq-normalized-yamabe}) with the initial functions $u, u_{i}$ respectively
exists up to $T.$
Hereafter, we fix such a positive existence time $T > 0$ (independent of $i$). 

From the assumption $(1)$ in Main Theorem \ref{main}, by taking $i$ large enough, we can assume that 
\[
\frac{1}{2}\, \mathrm{Vol}(M, g) \le \mathrm{Vol}(M, g_{i}) \le 2\, \mathrm{Vol}(M, g)~~\mathrm{for~all}~i,
\]
where $\mathrm{Vol}(M, h)$ is the volume of $M$ measured by the Riemannian metric $h.$
Moreover, since the normalized Yamabe flow preserves the volume, we obtain that
\[
\frac{1}{2}\, \mathrm{Vol}(M, g) \le \mathrm{Vol}(M, g_{i}(t)) \le 2\, \mathrm{Vol}(M, g)
\]
for all $i$ and $t \in [0, T]$ as well.
\begin{rema}
\label{rema-volume}
Since $g_{i} = u_{i}^{\frac{4}{n-2}} g_{0}$ and $g = u^{\frac{4}{n-2}} g_{0}$ respectively, we have respectively
\[
\mathrm{Vol}(M, g_{i}) = \int_{M} u_{i}^{\frac{2n}{n-2}}\, d\mathrm{vol}_{g_{0}}~\mathrm{and}~
\mathrm{Vol}(M, g) = \int_{M} u^{\frac{2n}{n-2}}\, d\mathrm{vol}_{g_{0}}.
\]
Hence the above volume estimates also hold under the weaker assumption $(1)'$ in Main Theorem \ref{main}.
\end{rema}
We will establish the convergence of $u_{i}(t)$ to $u(t)$ on $M \times [0, T]$ below.
We shall do this in two cases: $Y(M, g_{0}) \le 0$ and $Y(M, g_{0}) > 0.$

\medskip
\noindent
\underline{\textbf{Yamabe non-positive case (i.e., $Y(M, g_{0}) \le 0$)}}:
As stated in {\cite[Section~3]{ye1994global}}, since $Y(M, g_{0}) < 0~(\mathrm{resp.}~\le 0),$
we can choose a background metric $\tilde{g}_{0} \in [g_{0}]$ such that $R(\tilde{g}_{0}) < 0~(\mathrm{resp.}~\le 0)$ on $M.$
Set $\tilde{g}_{0} =: \tilde{u}^{-\frac{4}{n-2}} g_{0},$
then $u \cdot \tilde{u}$ and $u_{i} \cdot \tilde{u}$ satisfy the assumption $(1)$ in Main Theorem \ref{main}.
Thus, without loss of generality, we can assume that $R(g_{0}) < 0~(\mathrm{resp.}~\le 0)$ on $M.$
From the maximum principle {\cite[(3.1)]{ye1994global}}, we obtain that for all $t \in (0, T),$
\[
\frac{d}{dt} u_{i, \mathrm{min}}^{\frac{n+2}{n-2}}(t) \ge \frac{n-2}{4(n-1)} \min_{M} \left| R(g_{0}) \right|\, u_{i, \mathrm{min}}(t) + \frac{n-2}{4(n-1)} r(g_{i}(t))\, u_{i, \mathrm{min}}^{\frac{n+2}{n-2}}(t),
\] 
where $u_{i, \mathrm{min}}(t) := \min_{M} u_{i}(\cdot, t).$
By the definition of $Y(M, g_{0})$ and the above volume estimate (see also Remark \ref{rema-volume}), we have
$r(g_{i}(t)) \ge Y(M, g_{0}) \mathrm{Vol}(M, g_{i}(t))^{-\frac{2}{n}} \ge \frac{1}{2} Y(M, g_{0}) \mathrm{Vol}(M, g)^{-\frac{2}{n}}.$
Thus we get
\[
\frac{d}{dt} u_{i, \mathrm{min}}^{\frac{n+2}{n-2}}(t) \ge \frac{n-2}{8(n-1)} Y(M, g_{0}) \mathrm{Vol}(M, g)^{-\frac{2}{n}}\, u_{i, \mathrm{min}}^{\frac{n+2}{n-2}}(t)
\]
for all $t \in (0,T).$
Hence, from this estimate and the mean value theorem (cf. Remark \ref{rema-prop2.1}), we obtain that
\[
u_{i, \mathrm{min}}^{\frac{n+2}{n-2}}(t) \ge \exp \left( \frac{n-2}{8(n-1)} Y(M, g_{0}) \mathrm{Vol}(M, g)^{-\frac{2}{n}}\, t \right) \cdot u_{i, \mathrm{min}}^{\frac{n+2}{n-2}}(0)
\]
for all $t \in [0, T].$
On the other hand, the maximum principle {\cite[(3.3)]{ye1994global}} implies that
\[
\frac{d}{dt} u_{i, \mathrm{max}}^{\frac{n+2}{n-2}}(t) \le -\frac{n-2}{4(n-1)} \left( \min_{M} R(g_{0}) \right)\, u_{i, \mathrm{max}}(t) + \frac{n-2}{4(n-1)} r(g_{i}(t))\, u_{i, \mathrm{max}}^{\frac{n+2}{n-2}}(t),
\]
where $u_{i, \mathrm{max}} (t) := \max_{M} u_{i}(\cdot, t).$
This implies that
\[
\frac{d}{dt} u_{i, \mathrm{max}}^{\frac{4}{n-2}} \le - \frac{n-2}{(n-1)(n+2)} \left( \min_{M} R(g_{0}) \right) + \frac{n-2}{(n-1)(n+2)} r(g_{i}(t)) u^{\frac{4}{n-2}}_{i, \mathrm{max}}(t).
\]
Moreover, since $t \mapsto r(g_{i}(t))$ is decreasing, we have
\[
\frac{d}{dt} u_{i, \mathrm{max}}^{\frac{4}{n-2}} \le - \frac{n-2}{(n-1)(n+2)} \left( \min_{M} R(g_{0}) \right) + \frac{n-2}{(n-1)(n+2)} r(g_{i}(0)) u^{\frac{4}{n-2}}_{i, \mathrm{max}}(t).
\]
By Gronwall's inequality (Proposition \ref{prop-app} in Section \ref{appendix}), we obtain that
\[
u_{i, \mathrm{max}}^{\frac{4}{n-2}}(t) \le \alpha(t) + \int^{t}_{0} \alpha(s) \beta\,\exp \left( \beta (t-s) \right)\, ds~~\mathrm{for~all}~t \in [0,T],
\]
where
\[
\alpha(t) := u_{i, \mathrm{max}}(0)^{\frac{4}{n-2}} - \frac{n-2}{(n-1)(n+2)} \left( \min_{M} R(g_{0}) \right) t
\]
and
\[
\beta := \frac{n-2}{(n-1)(n+2)} \max \left\{ r(g_{i}(0)), 0 \right\}.
\]
From the assumption $(2)$ in Main Theorem \ref{main} and the lower volume estimate (Remark \ref{rema-volume}),
\[
\beta \le \frac{2(n-2)}{(n-1)(n+2)} \mathrm{Vol}(M, g)^{-1} \cdot \max \{ \kappa, 0 \} := \tilde{\beta}.
\]
Hence we obtain that
\[
u_{i, \mathrm{max}}^{\frac{4}{n-2}}(t) \le \alpha(t) + \int^{t}_{0} \alpha(s) \tilde{\beta}\, \exp \left( \tilde{\beta} (t-s) \right)\, ds~~\mathrm{for~all}~t \in [0,T].
\] 
Therefore there is a positive constant $C$ depending only on $\kappa, u, g_{0}$ and $M$
such that
\[
0 < C^{-1} \le u_{i}(x, t) \le C~~\mathrm{for~all}~(x, t) \in M \times [0,T]. 
\]
Unless otherwise stated, we will denote positive constants depending only on the given data ($\kappa, u, g_{0}$ and $M$) as the same symbol $C.$
Then, from Krylov-Safonov estimate (\cite{krylov1981certain}, see also {\cite[Theorem~12]{picard2019notes}} and {\cite[Proposition~4.2]{caldeira2021normalized}}), we obtain the following $C^{\alpha}$-estimate for some $\alpha \in (0, 1):$
\[
|| u_{i}(\cdot, \cdot) ||_{C^{\alpha/2, \alpha} \left (M \times [t_{0}, T] \right)} \le C~~\mathrm{for~all}~t_{0} \in (0,T)~\mathrm{and}~i.
\]
Since $u_{i}(t) \rightarrow u_{i}(0)$ in $W^{2,p}(M, g_{0})$ for all $p > n/2,$
by the Sobolev embedding: $W^{2,p} \hookrightarrow C^{0, \alpha}$ for sufficiently large $p >> n/2,$
we obtain that
$u_{i}(t) \rightarrow u_{i}(0)$ in $C^{0, \alpha}(M, g_{0}).$
Hence, as $t_{0} \rightarrow 0,$ we obtain that
\[
|| u_{i}(\cdot, \cdot) ||_{C^{\alpha/2, \alpha} \left (M \times [0, T] \right)} \le C.
\]
From this estimate, the regularity assumption of Main Theorem \ref{main} and the interior parabolic Schauder estimate ({\cite[Proposition~4.2.]{caldeira2021normalized}}), we obtain that
\[
|| u_{i}(\cdot, \cdot) ||_{C^{(2+\alpha')/2, 2 + \alpha'} \left( M \times [t, T] \right)} \le C~~\mathrm{for~all}~i~\mathrm{and}~t \in (0, T),
\]
where $\alpha' = \frac{n-2}{n+2} \alpha.$
Note here that we have used the fact that $t \mapsto r(g(t))$ is decreasing and the lower volume estimate (Remark \ref{rema-volume}) when we applied the Schauder estimate. 
From these estimate, we can choose a subsequence $(u_{i_{k}}(\cdot, t))_{t \in [0, T]}$ such that 
$u_{i_{k}}$ converges to some $C^{2}$ function $\tilde{u}$ on $M \times (0,T]$ in the locally uniformly $C^{2}$-sense in $M \times (0,T]$ and $\tilde{u}$ satisfies the normalized Yamabe flow equation (\ref{eq-normalized-yamabe}).
Moreover, from the above $C^{\alpha}$-estimate and the standard diagonal argument, we can choose the subsequence $u_{i_{k}}$ so that the limiting function $\tilde{u}$ is a $C^{2}$ function in $M \times (0,T]$ and a $C^{\beta}$ function on $M \times [0,T]$ for some $\beta < \alpha$, and $u_{i_{k}}$ converges to $\tilde{u}$ in the $C^{\beta}$-sense on $M \times [0,T]$.
In particular, since $u_{i_{k}}(\cdot, 0) \rightarrow u(\cdot)$ as $k \rightarrow \infty$
(from the assumption $(1)$ in Main Theorem \ref{main}),
$\tilde{u}(\cdot, 0) \equiv u(\cdot).$
Then, from the uniqueness of the normalized Yamabe flow ({\cite[Corollary~3.3]{caldeira2021normalized}}), we finally obtain that
$\tilde{u}(t) \equiv u(t)$ for all $t \in [0,T].$ 
\begin{rema}\label{rema-uniqueness}
  Our situation here and the one in {\cite[Corollary~3.3]{caldeira2021normalized}} are different.
  However, one can still apply the same arguments in {\cite[Corollary~3.3]{caldeira2021normalized}} to our setting and can obtain the uniqueness since $\tilde{u}(\cdot, 0) \equiv u(\cdot)$ and $\tilde{u},~u$ both satisfy the unnormalized Yamabe flow equation (\ref{eq-normalized-yamabe}) in $M \times (0, T]$. 
\end{rema}

\medskip
\noindent
\underline{\textbf{Yamabe positive case (i.e., $Y(M, g_{0}) > 0$)}}:
As done in the previous case, we can assume that $R(g_{0}) > 0$ without loss of generality.
Since the assumption $(1)'$ of Main Theorem \ref{main} is weaker than $(1),$
it is enough to prove the theorem under the assumption $(1)'.$ 
Then we can firstly obtain the following.
\begin{prop}[{\cite[Proposition~2.1.]{brendle2005convergence}}]
\label{prop-pres}
The scalar curvature of the metric $g_{i}(t)$ satisfies
\[
\inf_{M} R(g_{i}(t)) \ge \min \left\{ \delta, 0 \right\}~~\mathrm{for~all}~t \ge 0.
\]
\end{prop}
\begin{proof}
Since $g_{i}(t)$ satisfies the normalized Yamabe flow (\ref{eq-normalized-yamabe}) in $M \times (0,T),$ from {\cite[Proposition 2.1]{brendle2005convergence}}, 
\[
\inf_{M} R(g_{i}(t)) \ge \min \left\{ \delta, 0 \right\}~~\mathrm{for~all}~t \in (0, T).
\]
Then, the assertion follows from the $C^{0}$-continuity: $R(g_{i}(t)) \rightarrow R(g_{i})$ stated in Remark \ref{rema-prop2.1}.
\end{proof}
Set
\[
\sigma := \max \left\{ 1-\delta, 1 \right\},
\]
then $R(g_{i}(t)) + \sigma \ge 1$ for all $t \ge 0.$
Moreover, Brendle {\cite[Proposition~2.4]{brendle2005convergence}} also gave upper and lower bounds of $u_{i}(t).$
\begin{prop}[{\cite[Proposition~2.4]{brendle2005convergence}}]
There are positive constants $C$ and $c$ depending only on
$T, \delta, \kappa, g$ and $M$ such that 
\[
\sup_{M} u_{i}(t) \le C
\]
and
\[
\inf_{M} u_{i}(t) \ge c
\]
for all $t \in [0, T].$
\end{prop} 
\begin{proof}
The following proof mostly follow the argument of Brendle {\cite[Proposition~2.4]{brendle2005convergence}}.
The function $u_{i}(t)$ ($t \in (0, T)$) satisfies 
\[
\begin{split}
\frac{\partial}{\partial t} u_{i}(t) &= -\frac{n-2}{4} \left( R(g_{i}(t)) - r(g_{i}(t)) \right) \\
&\le \frac{n-2}{4} (r(g_{i}(0)) + \sigma) \\
&\le \frac{n-2}{4} \left( \frac{1}{2} \mathrm{Vol}(M, g)^{-1} \kappa + \sigma \right).
\end{split}
\]
We have used the assumption $(2)$ of Main Theorem \ref{main} and the volume lower estimate (Remark \ref{rema-volume}) in the last inequality.
From this estimate, the continuity of $u_{i}(t)$ as $t \rightarrow 0$ (see Remark \ref{rema-prop2.1}) and the mean value theorem, we obtain that
\[
\sup_{M} u_{i}(t) \le C~~\mathrm{for~all}~t \in [0, T].
\]
Moreover from this estimate, we follow the same argument in {\cite[Proposition~2.4]{brendle2005convergence}} and obtain the uniformly lower bound on $M \times [t_{0}, T]$ for all $t_{0} \in (0, T).$
Thus, we can obtain the lower estimate on the whole space-time $M \times [0,T]$ by the continuity of $u_{i}(t)$ as $t \rightarrow 0$ (see Remark \ref{rema-prop2.1}).
\end{proof}
\begin{rema}
  In order to obtain the uniform lower bound of $\inf_{M} u_{i}$ here, we have specifically used the fact that the volume $\mathrm{Vol}(M, u_{i}(t)^{\frac{4}{n-2}} g_{0})$ is bounded from below by some positive constant independent of time. Especially in this case, it is constant in time.
\end{rema}
Hence, from the same argument in the previous case, we can obtain that there is a function $\tilde{u}$, which is $C^{2}$ in $M \times (0,T]$ and $C^{\beta}$ for some $\beta < \alpha$ on $M \times [0,T]$ such that $u_{i}$ subconverges to $\tilde{u}$ in the locally uniformly $C^{2}$-sense in $M \times (0,T]$ and $C^{\beta}$-sense on $M \times [0,T]$.
From the assumption $(1)'$ in Main Theorem \ref{main}, as $i \rightarrow \infty,$ we obtain that 
\[
\int_{M} \left| \tilde{u}(\cdot, 0) - u \right|^{\frac{2n}{n-2}} d\mathrm{vol}_{g_{0}} = 0.
\]
Hence we obtain that $\tilde{u}(\cdot, 0) \equiv u.$
Indeed, if not, there is a point $p \in M$ such that $\tilde{u}(\cdot, 0)(p) - u(p) \neq 0.$
Without loss of generality, we may assume that $\tilde{u}(\cdot, 0)(p) - u(p) := \varepsilon > 0.$
Since $\tilde{u}(\cdot, 0) - u$ is continuous on $M,$ there are sufficiently small open neighborhoods $U, V$ of $p$ such that
$\tilde{u}(\cdot, 0) - u \ge \varepsilon / 2$ on $V$ and the closure of $V$ is contained in $U.$
Let $\psi \in C^{\infty}(M)$ be a smooth cut-off function such that
$0 \le \psi \le 1,$
$\psi \equiv 1$ on $V$ and  $\psi \equiv 0$ on $M \setminus U.$
Then, 
\[
\begin{split}
0 &= \int_{M} \left| \tilde{u}(\cdot, 0) - u \right|^{\frac{2n}{n-2}} d\mathrm{vol}_{g_{0}} \\
&= \int_{M} \left| \tilde{u}(\cdot, 0) - u \right|^{\frac{2n}{n-2}} \left( \left( 1-\psi \right) + \psi \right)\, d\mathrm{vol}_{g_{0}} \\
&\ge \int_{M} \left| \tilde{u}(\cdot, 0) - u \right|^{\frac{2n}{n-2}} \cdot \psi\, d\mathrm{vol}_{g_{0}} \\
&\ge \left( \frac{\varepsilon}{2} \right)^{\frac{2n}{n-2}} \cdot \mathrm{Vol}(V, g_{0}) > 0
\end{split}
\]
This is a contradiction.

Therefore, from the uniqueness of the normalized Yamabe flow ({\cite[Corollary~3.3]{caldeira2021normalized}} and see Remark \ref{rema-uniqueness}), we obtain that
$\tilde{u}(t) \equiv u(t)$ for all $t \in [0,T].$ 
 
\begin{rema}
One can follow the arguments of Brendle {\cite[Section~2]{brendle2005convergence}}
and we can also obtain $C^{\alpha}$-estimates ({\cite[Proposition~2.6]{brendle2005convergence}}).
However the constant $C$ in {\cite[Lemma~2.5]{brendle2005convergence}} depends not only 
$T, \delta, \kappa, g_{0}$ and $n$ but also 
\[
\int_{M} |R(g_{i}(0)) - r(g_{i}(0))|^{\frac{n^{2}}{2(n-2)}}\, d\mathrm{vol}_{g_{i}(0)},
\]
hence we need an additional information about such an $L^{\frac{n^{2}}{2(n-2)}}$-quantity.
Therefore we have used the Krylov--Safonov estimate \cite{krylov1981certain} instead here.
\end{rema}

\bigskip
Here, we give a proof of Main Theorem \ref{main}. 
(cf. \cite{bamler2016proof} and the author's preprint, arXiv:2208.01865v14)
\begin{proof}[Proof of Main Theorem \ref{main}]
Since $t \mapsto r(g_{i}(t))$ is decreasing and the normalized Yamabe flow (\ref{eq-normalized-yamabe}) preserves the volume, we have
\[
\kappa \ge \int_{M} R(g_{i}(0))\, d\mathrm{vol}_{g_{i}(0)} \ge \int_{M} R(g_{i}(t))\, d\mathrm{vol}_{g_{i}(t)}
\]
for all $t \in [0, T]$ and $i \in \mathbb{N}.$
From the precompactness of the normalized Yamabe flow stated above, (after taking a subsequence) as $i \rightarrow \infty,$ we have
\[
\int_{M} R(g(t))\, d\mathrm{vol}_{g(t)} \le \kappa~~\mathrm{for ~all}~t \in (0, T].
\]
Then, from the $C^{0}$-continuity of $g(t)$ and $R(g(t))$ as $t \rightarrow 0$ (see Remark \ref{rema-prop2.1}),
\[
\int_{M} R(g)\, d\mathrm{vol}_{g} = \int_{M} R(g(0))\, d\mathrm{vol}_{g(0)} 
= \lim_{t \rightarrow 0} \int_{M} R(g(t))\, d\mathrm{vol}_{g(t)} \le \kappa.
\]
\end{proof}

\begin{rema}
\begin{itemize}
\item[(1)] Since 
\[
\begin{split}
\int_{M} &R(g_{i}(t))\, d\mathrm{vol}_{g_{i}(t)} \\
&= \int_{M} - u_{i}(t)^{-\frac{n+2}{n-2}} \left( \frac{4(n-1)}{n-2} \Delta_{g_{0}} u_{i}(t) - R(g_{0}) u_{i}(t) \right) \cdot u_{i}(t)^{\frac{2n}{n-2}}\, d\mathrm{vol}_{g_{0}} \\
&=  \int_{M} \left( \frac{4(n-1)}{n-2} |\nabla u_{i}(t)|^{2} + R(g_{0}) u_{i}(t)^{2} \right)\, d\mathrm{vol}_{g_{0}},
\end{split}
\]
we have actually used only the fact that $g_{i}(t)$ converges to $g(t)$ in the locally uniformly $C^{1}$-sense in $M \times (0,T]$
to verify that
$\int_{M} R(g(t))\, d\mathrm{vol}_{g(t)} \le \kappa$ for all $t \in (0, T]$.

\item[(2)] Assume the same as in Main Theorem \ref{main} with $\delta \le 0$ except $(2),$ and assume that $R(g_{0}) \le 0.$
Then we can prove that
\[
R(g) \ge \delta~~\mathrm{on}~M.
\]
In other words, we can give a partial proof of the Gromov's result \cite{gromov2014dirac} for this case.
Indeed, in our proof of Main Theorem \ref{main} above, $g_{i}(t)$ converges to $g(t)$ on $M \times (0, T]$ in the uniformly $C^{2}$-sense and the lower bound of the scalar curvature
$R(g_{i}(t))$ was preserved under the flow for each $i$, which follows from Proposition \ref{prop-pres}.
Therefore, by the same argument of the proof of Main Theorem \ref{main}, we can prove this claim (see also \cite{bamler2016proof} and the author's preprint, arXiv:2208.01865v14).

\item[(3)] In the argument in the Yamabe non-positive case,
we cannot obtain in general that 
the sequence of the solution $u_{i}$ of the unnormalized Yamabe flow can subconverge to  the limit $u(t)$ up to $C^{1}$-sense on the whole space-time $M \times [0,T].$
Indeed, the sequence of metrics constructed in Example 3.3 of the author's preprint: arXiv:2208.01865v14 has a uniform upper bound of the total scalar curvature, however it does not converge to the limiting metric in the $C^{1}$-sense.
\end{itemize}
\end{rema}

\section{Proof of Main Theorem \ref{main2}}

\,\,\,\,\,\,\, As done in the previous section, we can assume that $R(g_{0}) < 0$ (resp. $\le 0$) when $Y(M, g_{0}) < 0$ (resp. $Y(M, g_{0}) \le 0$).
Then, from the proof of Proposition \ref{prop-short} and the argument in the previous section (Yamabe non-positive case), we can also obtain the long-time solution of the unnormalized Yamabe flow:
\begin{equation}\label{eq-yamabe-flow}
\begin{split}
  \frac{\partial}{\partial t} g_{i}(t) &= -R(g_{i}(t))\, g_{i}(t)~~~~\left( t \in [0, \infty) \right)~~\mathrm{and} \\
  \frac{\partial}{\partial t} g(t) &= -R(g(t))\, g(t)~~~~\left( t \in [0, \infty) \right) 
\end{split}
\end{equation}
respectively with $g_{i}(0) = g_{i}$ and $g(0) = g$ respectively.
(Compared with the situation in Section \ref{section-proof}, the term of the mean scalar curvature $r(g_{i}(t))$ does not appear.)
In order to prove that the unique short-time solution obtained by Proposition \ref{prop-short} is actually long-time solution, it is sufficient to prove that the $C^{0}$-norm of the obtained solution is uniformly bounded from above and below in the space-time. (cf. \cite{brendle2005convergence, carron2021convergence})
More precisely, we can obtain such uniform bounds in the same way as we did in ``Yamabe non-positive case'' of the proof of Main Theorem \ref{main}.
However, the volume is not invariant under the flow in this case, hence we use the following volume estimate (especially from below) instead.
\begin{lemm}[Volume estimates along the unnormalized Yamabe flow]
Under the assumption of Main Theorem \ref{main2}, along the unnormalized Yamabe flow $g(t)$ (\ref{eq-yamabe-flow}),
{\small \[
\mathrm{Vol}(M, g(0)) -\frac{n \kappa}{2} t
\le \mathrm{Vol}(M, g(t)) \le \left( -t Y(M, g_{0}) + \mathrm{Vol}(M, g(0))^{\frac{2}{n}} \right)^{\frac{n}{2}}.
\]}
\end{lemm}
\begin{proof}
Along the flow, the volume evolves as 
\[
\frac{\partial}{\partial t} \mathrm{Vol}(M, g(t)) = -\frac{n}{2} \int_{M}R(g(t))\, d\mathrm{vol}_{g(t)}.
\]
On the other hand, one can easily check that $t \mapsto \int_{M} R(g(t))\, d\mathrm{vol}_{g(t)}$ is decreasing (see the proof of Main Theorem \ref{main2} below), hence
\[
\begin{split}
-\frac{n}{2} \int_{M}R(g(t))\, d\mathrm{vol}_{g(t)} &\ge -\frac{n}{2} \int_{M} R(g(0))\, d\mathrm{vol}_{g(0)} \\
&\ge -\frac{n}{2} \kappa.
\end{split}
\]
Here, we have used the assumption $(c)$ of Main Theorem \ref{main2} in the last inequality.
On the other hand, from the definition of $Y(M, g_{0}),$ we have
\[
-\frac{n}{2} \int_{M}R(g(t))\, d\mathrm{vol}_{g(t)} \le -\frac{n}{2} Y(M, g_{0}) \mathrm{Vol}(M, g(t))^{\frac{n-2}{n}}.
\]
Therefore, from these estimates, the continuity of $u(t)$ as $t \rightarrow 0$ and the mean value theorem, we obtain the desired assertion.
\end{proof}
Thus, we can take a positive time $T > 0$ (independent of $i$) and consider the unnormalized Yamabe flows $(g_{i}(t))_{t \in [0,T]}$ and $(g(t))_{t \in [0,T]}.$
We set $g_{i}(t) =: u_{i}(t)^{\frac{4}{n-2}} g_{0}$ and $g(t) =: u(t)^{\frac{4}{n-2}} g_{0},$
then $u_{i}(0) = u_{i}$ and $u(0) = u.$
In order to prove Main Theorem \ref{main2}, we also require the following $L^{1}$-estimate.
\begin{lemm}[cf. {\cite[Theorem 2.2]{takahashi2022infinite}}]
\label{lemm-esti}
Under the above settings, for any nonnegative smooth function $\psi \in C^{\infty}(M),$ we have
\begin{equation}
\label{eq-L1}
\begin{split}
&\left( \int_{M} \psi |u(t) - u_{i}(t)|\, d\mathrm{vol}_{g_{0}} \right)^{\frac{4}{n+2}} \\
&\le \left( \int_{M} \psi |u(0) - u_{i}(0)|\, d\mathrm{vol}_{g_{0}} \right)^{\frac{4}{n+2}} \\
&~~~+ \left( \frac{(n-1)(n+2)}{n-2} \left( 2 C[\psi] \right)^{\frac{4}{n+2}} + \frac{n+2}{4} \left( \int_{M} \psi\, d\mathrm{vol}_{g_{0}} \right)^{\frac{4}{n+2}} \right)\, t,
\end{split}
\end{equation}
where
\[
C[\psi] := \int_{M} |\Delta_{g_{0}} \psi|^{\frac{n+2}{4}} \psi^{-\frac{n-2}{4}}\, d\mathrm{vol}_{g_{0}}.
\]
\end{lemm}
\begin{proof}
Since $u(t)$ and $u_{i}(t)$ satisfy the unnormalized Yamabe flow equation (\ref{eq-yamabe-flow}) and are positive values, we can obtain the following estimate (see {\cite[Proof of Theorem 2.2.]{takahashi2022infinite}}):
For any nonnegative smooth function $\psi \in C^{\infty}(M),$
\[
\begin{split}
\frac{d}{dt} \int_{M} \psi |u(t) - &u_{i}(t)|\, d\mathrm{vol}_{g_{0}} \\
&\le \frac{(n-1)(n+2)}{n-2} \int_{M} \Delta_{g_{0}} \psi \left| u(t)^{\frac{n-2}{n+2}} - u_{i}(t)^{\frac{n-2}{n+2}} \right|\, d\mathrm{vol}_{g_{0}} \\
&~~~+ \frac{n+2}{4} \int_{M} \psi \left| R(g_{0}) \right| \left| u(t)^{\frac{n-2}{n+2}} - u_{i}(t)^{\frac{n-2}{n+2}} \right|\, d\mathrm{vol}_{g_{0}}.
\end{split}
\]
then, from the same argument in {\cite[Proof of Theorem 2.2.]{takahashi2022infinite}},
we obtain that
\[
\begin{split}
\frac{d}{dt} \int_{M} &\psi |u(t) - u_{i}(t)|\, d\mathrm{vol}_{g_{0}} \\
&\le \frac{(n-1)(n+2)}{n-2} \left( 2 C[\psi] \right)^{\frac{4}{n+2}} \left( \int_{M} \psi |u(t) - u_{i}(t)|\, d\mathrm{vol}_{g_{0}} \right)^{\frac{n-2}{n+2}} \\
&+ \frac{n+2}{4} \left( \int_{M} \psi\, d\mathrm{vol}_{g_{0}} \right)^{\frac{4}{n+2}} \left( \int_{M} \psi |u(t) - u_{i}(t)|\, d\mathrm{vol}_{g_{0}} \right)^{\frac{n-2}{n+2}},
\end{split}
\]
where
\[
C[\psi] := \int_{M} |\Delta_{g_{0}} \psi|^{\frac{n+2}{4}} \psi^{-\frac{n-2}{4}}\, d\mathrm{vol}_{g_{0}}.
\]
From this, we obtain that 
\[
\begin{split}
\frac{d}{dt} &\left( \int_{M} \psi |u(t) - u_{i}(t)|\, d\mathrm{vol}_{g_{0}} \right)^{\frac{4}{n+2}} \\
&\le \frac{(n-1)(n+2)}{n-2} \left( 2 C[\psi] \right)^{\frac{4}{n+2}} 
+ \frac{n+2}{4} \left( \int_{M} \psi\, d\mathrm{vol}_{g_{0}} \right)^{\frac{4}{n+2}}.
\end{split}
\]
Integrating both sides of the previous inequality in time, we finally obtain that 
\[
\begin{split}
&\left( \int_{M} \psi |u(t) - u_{i}(t)|\, d\mathrm{vol}_{g_{0}} \right)^{\frac{4}{n+2}} \\
&\le \left( \int_{M} \psi |u(0) - u_{i}(0)|\, d\mathrm{vol}_{g_{0}} \right)^{\frac{4}{n+2}} \\
&~~~+ \left( \frac{(n-1)(n+2)}{n-2} \left( 2 C[\psi] \right)^{\frac{4}{n+2}} + \frac{n+2}{4} \left( \int_{M} \psi\, d\mathrm{vol}_{g_{0}} \right)^{\frac{4}{n+2}} \right)\, t.
\end{split}
\]
\end{proof}
By the assumption $(a)$ and the same argument in the previous section (Yamabe non-positive case),
we also obtain the following $C^{\alpha}$-estimate.
\begin{lemm}
Under the condition of Main Theorem \ref{main2}, there is a positive constant $C$ depending only on $M, u,,g_{0}, C_{0}$ and a constant $\kappa$
such that for some $\alpha \in (0,1),$
\[
|u(x_{1}, t_{1}) - u(x_{2}, t_{2})| \le C \left( |t_{1} - t_{2}|^{\frac{\alpha}{2}} + d_{g_{0}}(x_{1}, x_{2})^{\alpha} \right)
\]
and 
\[
|u_{i}(x_{1}, t_{1}) - u_{i}(x_{2}, t_{2})| \le C \left( |t_{1} - t_{2}|^{\frac{\alpha}{2}} + d_{g_{0}}(x_{1}, x_{2})^{\alpha} \right)
\]
for all $i.$
\end{lemm}
Combining these, we can show that $u_{i}$ actually converges to $u$ $C^{0}$-uniformly on $M$ as $i \rightarrow \infty.$
\begin{lemm}
\label{lemm-key}
$u_{i} \rightarrow u$ uniformly on $M$ as $i \rightarrow \infty.$
\end{lemm} 
\begin{proof}
We prove this lemma by contradiction.
Suppose that there is $\varepsilon > 0$ and
a sequence $\{ x_{i} \}$ of points in $M$ such that
\[
|u(x_{i}) - u_{i}(x_{i})| \ge \varepsilon.
\]
Fix a sufficiently small positive constant $0 < r_{0} << \min \{ \mathrm{inj}(M, g_{0}), \varepsilon \},$ where $\mathrm{inj}(M, g_{0})$ is the injectivity radius of $(M, g_{0}).$
In particular, 
\[
|u(x_{i}) - u_{i}(x_{i})| \ge \varepsilon > r_{0} > 0.
\]
We will denote some positive constants depending only on $M, g_{0}, u, \kappa$ and $C_{0}$ as the same symbol $C.$
We take a nonnegative test function $\psi \in C^{\infty}(M)$ such that 
\begin{itemize}
\item $\psi$ has compact support contained in the interior of 
\[
\{ x \in M |~d_{g_{0}}(x_{i}, x) \le r_{0} \},
\]
\item $0 \le \psi \le 1$ on $M.$
\end{itemize}
From the above $C^{\alpha}$-estimate and the volume comparison,
the left-hand side of (\ref{eq-L1}) (in the case of $t = 0$) in Lemma \ref{lemm-esti} is bounded from below by
\[
C r_{0}^{\left( 1 + \frac{n}{\alpha} \right) \frac{4}{n+2}}.
\]
On the other hand, the right-hand side of (\ref{eq-L1}) ($t = 0$) is bounded from above by 
\[
\left( \int_{M} |u(0) - u_{i}(0)|\, d\mathrm{vol}_{g_{0}} \right)^{\frac{4}{n+2}}.
\]
However, from the assumption $(b),$ we can take $i$ sufficiently large such that
\[
\left( \int_{M} |u(0) - u_{i}(0)|\, d\mathrm{vol}_{g_{0}} \right)^{\frac{4}{n+2}} < \frac{C}{2} r_{0}^{\left( 1 + \frac{n}{\alpha} \right) \frac{4}{n+2}}.
\]
This is a contradiction.
\end{proof}

\begin{proof}[Proof of Main Theorem \ref{main2}]
Along the unnormalized Yamabe flow $(g(t))_{t \in [0,T)},$ the scalar curvature evolves as 
\[
\frac{\partial}{\partial t} R(g(t)) = (n-1) \Delta_{g(t)} R(g(t)) + R(g(t))^{2}.
\]
Thus, we obtain that for all $t \in (0,T),$
\[
\begin{split}
\frac{d}{dt} &\left( \int_{M} R(g(t))\, d\mathrm{vol}_{g(t)} \right) \\
&= \int_{M} \left( (n-1)\Delta_{g(t)} R(g(t)) + R(g(t))^{2} - \frac{n}{2} R(g(t))^{2} \right)\, d\mathrm{vol}_{g(t)} \\
&= -\frac{n-2}{2}\int_{M} R(g(t))^{2}\, d\mathrm{vol}_{g(t)} \\
&\le 0.
\end{split}
\]
Hence the function $t \mapsto \int_{M} R(g(t))\, d\mathrm{vol}_{g(t)}$ is decreasing for all $t \in [0,T]$ because $R(g(t)) \overset{C^{0}}{\longrightarrow} R(g(0))$ (see Remark \ref{rema-prop2.1}).
Therefore, from Lemma \ref{lemm-key}, we can follow the same argument in the proof of Main Theorem \ref{main} and prove Main Theorem \ref{main2} in the same manner.
\end{proof}

\begin{ques}
We have the following questions. 
\begin{itemize}
\item On an open manifold, is there example for which Main Theorems does not hold?
\item Is the condition ``$\int_{M} R(g)\, dm \le \kappa$'' for Riemannian manifolds with smooth measures $(M, g, dm := e^{-f} d\mathrm{vol})$ $C^{0}$-closed
in the intersection of a conformal class and the space of all Riemannain metrics with measures?  (cf. Main Theorem 2 of the author's preprint, arXiv:2208.01865v14)
\end{itemize}
\end{ques}

\section{Appendix}
\label{appendix}
We used the following fact in the proof of Main Theorem \ref{main} (in the Yamabe non-positive case).
\begin{prop}[Gronwall's inequality]
\label{prop-app}
Assume that there is a continuous function $\alpha(t)$ and a nonnegative continuous function $\beta(t)$ on $[0, T]$ such that a continuous function $u(t)$ satisfies the following inequality:
\[
u(t) \le \alpha(t) + \int_{0}^{t} \beta(s) u(s)\, ds~~~~\mathrm{for~all}~t \in [0, T].
\]
Then, for all $t \in [0, T],$
\[
u(t) \le \alpha(t) + \int_{0}^{t} \alpha(s) \beta(s) \exp \left( \int_{s}^{t} \beta(r)\, dr \right)\, ds.
\]
\end{prop}
\begin{proof}
Set 
\[
v(s) := \exp \left( -\int_{0}^{s} \beta(r)\, dr \right) \cdot \int_{0}^{s} \beta(r) u(r)\, dr
\]
for $s \in [0, T].$
Then, we can easily check that
\[
\frac{d}{ds} v(s) = \left( u(s) - \int_{0}^{s} \beta(r) u(r)\,dr \right) \beta(s) \exp \left( -\int_{0}^{s} \beta(r)\, dr \right),~~~s \in [0,T].
\]
Since $v(0) = 0,$ using the assumption and integrating the previous inequality, we obtain that
\[
v(t) \le \int_{0}^{t} \alpha(s) \beta(s) \exp \left( -\int_{0}^{s} \beta(r)\, dr \right)\, ds
\]
for all $t \in [0, T].$
Thus, from the definition of $v(t),$
\[
\begin{split}
\int_{0}^{t} \beta(s) u(s)\, ds &= \exp \left( \int_{0}^{t} \beta(r)\, dr \right) v(t) \\
&\le \int_{0}^{t} \alpha(s) \beta(s) \exp \left( \int_{0}^{t} \beta(r)\, dr - \int_{0}^{s} \beta(r)\, dr \right)\, ds \\
&= \int_{0}^{t} \alpha(s) \beta(s) \exp \left( \int_{s}^{t} \beta(r)\, dr \right)\, ds.
\end{split}
\]
Substituting this into the assumed inequality, we can obtain the desired assertion. 
\end{proof}

\bigskip
\noindent
\textit{E-mail adress}:~hamanaka1311558@gmail.com

\smallskip
\noindent
\textsc{Department of Mathematics, Graduate School of Science, Osaka University, Toyonaka, Osaka 560-0043, Japan}

\end{document}